OpenStudy (anonymous):

x^-3 = 64 . Show work please? I want to know how to work problems like this faster

OpenStudy (anonymous):

x^-3=64 1/x^3=4^3 1/4^3=x^3 (1/4)^3=x^3 so x=1/4

OpenStudy (anonymous):

1/x^3=64 1/64=x^3 (1/4)^3=x^3 x=1/4 Check: (1/4)^(-3)=64 4^3=64 True